# The stochastic gravitational wave background from cosmic superstrings

### v1

@@ Abstract
We study the stochastic gravitational wave background sourced by a network of cosmic superstrings and demonstrate that incorporating higher-mass string species, beyond the fundamental string, is crucial for accurately modeling the resulting gravitational wave spectrum across frequencies ranging from nanohertz to kilohertz. Using the multi-tension velocity-dependent one-scale model to evolve the cosmic superstring network, we perform several fits to the NANOGrav 15-year dataset and obtain expectation values for the fundamental string tension, string coupling and effective size of compact extra dimensions. We find that the cosmic superstring best-fits are comparable in likelihood to Supermassive Black Hole models, thought by many to be the leading candidate explanation of the signal. The implications of the best-fit spectra are discussed within the context of future gravitational wave experiments. We obtain expectation values for the fundamental string tension of $\log_{10}(G\mu_{1})=-11.5^{+0.3}_{-0.3}$($-11.6^{+0.2}_{-0.3}$) for gravitational waves originating from large cuspy (kinky) cosmic superstring loops and $\log_{10}(G\mu_{1})=-9.7^{+0.7}_{-0.7}$($-9.9^{+1.0}_{-0.5}$) for small cuspy (kinky) loops. We also place $2\sigma$ upper bounds on the string coupling, finding $g_{s}<0.65$ in all cases, and comment on the implication of our results for the effective size of the compact extra dimensions.
@@ I Introduction
Cosmic superstrings occur in certain string theory scenarios where fundamental strings (F-strings) and one-dimensional Dirichlet branes (D-strings) are stretched to cosmic scales by brane inflation Sarangi and Tye (2002); Jones et al. (2002); Majumdar and Christine-Davis (2002); Copeland et al. (2004). Unlike ordinary cosmic string networks, which generally consist of a single type of string with unit chopping probability when intersecting, cosmic superstrings have multiple different species (the F- and D-strings together with their bound states) and a reduced probability of intersecting. As they are string theory objects, the extra dimensions they experience along with their more complicated dynamics significantly increases the difficulty of modelling the evolution of these networks. Nevertheless, remarkable strides have been made towards a consistent picture of cosmic superstring network evolution by modifying methods used to study cosmic strings Avgoustidis and Copeland (2010); Avgoustidis and Shellard (2008); Pourtsidou et al. (2011); Marfatia and Zhou (2024); Correia and Martins (2022); Blanco-Pillado et al. (2021). While there remain uncertainties in the detailed behaviour and interactions of cosmic superstring networks, we believe that we are at a point where detailed predictions can be made about the spectrum of stochastic gravitational waves arising from them Sousa and Avelino (2016).
@@ II.1 String Dynamics
where $\mathcal{P}_{i}\equiv\mathcal{P}_{ii}$ is the microphysical self-interaction probability for strings of type $i$. The prefactor $\tilde{c}$ ensures that for ${\mathcal{P}}_{i}=1$ one recovers the standard VOS value $\tilde{c}$ Martins and Shellard (2002) for the loop chopping efficiency, while the power $1/3$ is supported Avgoustidis and Shellard (2006) by numerical simulations of Nambu-Goto strings with a suppressed intercommutation probability, in both matter and radiation dominated cosmologies222 Earlier simulations of Nambu-Goto strings evolving in Minkowski spacetime may suggest a slightly stronger dependence Sakellariadou and Vilenkin (1990); Dvali and Vilenkin (2004) of $\tilde{c}$ on the intercommuting probability $P$, namely $\tilde{c}\propto P^{1/2}$.. Note that going from $\tilde{d}_{ij}$ in equation (5) to $\tilde{d}^{k}_{ij}=\tilde{d}_{ij}S^{k}_{ij}$ in equations (2-3) involves a kinematic factor $S^{k}_{ij}$ Avgoustidis and Copeland (2010); Pourtsidou et al. (2011) (also see Copeland et al. (2007); Salmi et al. (2008); Copeland et al. (2008)) quantifying the relative probability that the collision proceeds via the additive or subtractive channel Tye et al. (2005), i.e. $(p,q)+(p^{\prime},q^{\prime})\longrightarrow(p\pm p^{\prime},q\pm q^{\prime})$, producing a heavier or lighter string segment, respectively.
@@ III.1 Importance of heavier strings
We start by confirming that, as shown in Sousa and Avelino (2016), it is crucial to model the gravitational wave contributions of strings heavier than the fundamental string, as they give rise to additional structure in the spectrum. Since the cutoff frequency of GW emission of a given string species, $i$, is given by [^5] Recall we have the loop size given in terms of the time of chopping by $l_{i}=\alpha_{i}t$, and the fundamental frequency of the GW it emits is given by $f_{i}=2/l_{i}$. $f_{\rm min,i}=2/(\alpha_{i}t_{0})$, and since $\alpha_{i}$ can vary between string species, the cut-off frequency becomes species-dependent. If $\alpha_{i}$ differs between species, this leads to regions of the frequency space where the GW emission has stopped for fundamental strings, but continues for heavier strings, as discussed earlier (cf. low frequency regions of Fig. 2 corresponding to small loops with $\alpha_{i}=2\Gamma G\mu_{i}$) for $g_{s}=0.04$ and $g_{s}=0.3$. The result is a multi-peaked spectrum at low frequencies – a clear signature of the GWs arising from a multi-tension network.
However, this is not the whole story. If $\alpha_{i}$ is the same for all species, we do not see multiple peaks but there can still be areas of the frequency space where heavier strings dominate over the fundamental string at low frequency (see leftmost plot of Fig. 3 corresponding to large loops with $\alpha_{i}=0.1$ for $g_{S}=0.04$). As loop birth times depend on the tension of the string species (and loop density is cut off when $t<t_{b}$), the integral in Eq. (16) can be smaller for the fundamental strings than for heavier strings, even though the smaller correlation lengths result in many more loops overall. It follows that GW contributions from heavier string species are strongest for low values of the string coupling (as these enhance the heavier string tension – see Eq. (1)) and in cases where $\alpha_{i}$ is species dependent (e.g. when loops are produced at gravitational back-reaction scales).
@@
Our results, as summarised in Figs. 2 and 3 are in agreement with reference Sousa and Avelino (2016), up to changes in the locations of the peaks caused by different choices/interpretations of the parameter $\alpha_{i}$, which in the case of reference Sousa and Avelino (2016) corresponds to $\alpha_{Li}$ in our notation. Thus, the physical loop sizes differ by a factor of the time normalised correlation length $\xi_{i}$. In particular, in our interpretation of the “large loop” value $\alpha_{i}=0.1$ the physical size of the loops is the same for all strings and so there are no multiple peaks in Fig. 3. Other, smaller, differences arise from our more accurate modelling of the cosmic superstring network achieved by (a) including decay processes of heavier strings down to the three lighest species (as discussed above), and (b) by taking into account the dependence of the VOS parameters $\tilde{c_{i}}$, $\tilde{d}^{k}_{ij}$ on the scaling velocities of each string species.
Figure 3: Plots showing the gravitational wave background spectra generated at different string coupling constants at string tension $G\mu_{1}=10^{-9}$ and loop size $\alpha_{i}=0.1$. The solid lines show the GWs from the 7-string model and the dashed lines from the 3-string model.
@@ III.2.2 Loop size, $\alpha_{i}$
One of the most important parameters in determining the shape and magnitude of the GW background spectrum is the size of string loops when they are chopped off from the string network. Motivated by the scaling behaviour of cosmic (super)string networks, it is commonly assumed that their size is a constant fraction of the string correlation length at the time of birth, so that $l_{i}(t_{b})=\alpha_{i}t_{b}$, where $\alpha_{i}$ is a constant. The exact value of $\alpha_{i}$ is unknown. Throughout the years, investigations into cosmic strings have shown loops produced at sizes comparable to the string thickness Bevis et al. (2008); Vincent et al. (1997a), sizes around the gravitational backreaction scale, $\alpha_{i}=2\Gamma G\mu_{i}$ Bennett and Bouchet (1990); Albrecht and Turok (1985); Vincent et al. (1997b); Allen and Shellard (1990), and at scales much closer to the characteristic length of the string network Olum and Vanchurin (2007); Martins and Shellard (2006); Ringeval et al. (2007); Vanchurin et al. (2005, 2006). The most recent large-scale numerical simulation into the production of cosmic string loops finds a double-peaked spectrum of loop production, where around 90% of network energy goes into loops at the gravitational backreaction scale and the remaining 10% into large loops with $\alpha_{i}\sim 0.1$ Blanco-Pillado et al. (2014). Note in that paper the size of loops is expressed as a fraction of the horizon size during the radiation era, i.e $\alpha_{h}$, which they define as $\alpha$. This differs from the definition used in this work by a factor of 2.
Given the lack of a clear consensus, we will look at GWs emitted at the two scales found by Blanco-Pillado et al. (2014) as these will allow us to look at both the large and small loop physics of GW emission. Note that it is not clear that results derived from cosmic string simulations can be directly applied to cosmic superstrings and $\alpha_{i}$ can, in principle, take any other value. Nevertheless, we would expect these scales to be representative of the range of possible values.
@@ III.2.4 Size of the extra dimensions and energy of the loops
As discussed in section II, the volume parameter $w$ is the ratio between the minimum volume and effective volume occupied by the fundamental string in the compact extra dimensions. A small value of $w$ corresponds to a larger volume of the compact dimensions being available to the strings, while $w=1$ corresponds to compactification at the string scale (or to highly warped extra dimensions) so that volume effects from the extra dimensions become negligible. Note that the effective volume is only a lower bound to the total volume of the extra compact dimensions. As more of the extra dimensions become traversable to the strings, the larger the probability of the strings “missing” each other in the extra dimensions when intersecting. This leads to a decrease in intercommuting probability, corresponding to an increase in the number of string loops and a boost to the GW signal. Since the heavier strings oscillate less in the extra dimensions, the effective volume they explore is much smaller than for the fundamental strings and the boost in GW magnitude is much smaller. Comparison spectra for $w=0.1$ and $w=1$ can be seen in Fig. 4. As varying $w$ has a strong impact on the resulting spectra and the value of $w$ is physically interesting, encoding information on the size/warping of the extra dimensions, we will treat this as a free parameter.
@@ IV Results and Discussion
Cuspy big loops $47.7$ $-11.5^{+0.3}_{-0.3}$ $<0.65$ $<0.91$
-12.0 0.14 0.02
Kinky big loops $47.0$ $-11.6^{+0.2}_{-0.3}$ $<0.65$ $<0.92$
-12.4 0.10 0.01
@@ IV.3 Comparison to NANOGrav fit
where $\Omega_{\rm CS}$ is the spectrum obtained from cosmic strings at a given tension and $P$ is treated as a constant free parameter [^6] The scaling $\Omega_{\rm SUPER}(f)\propto P^{-1}$ can be understood in terms of the VOS model as taking the loop chopping efficiency $\tilde{c}\propto P^{1/2}$ leading to a scaling $L\propto P^{1/2}$ in the correlation length Sakellariadou and Vilenkin (1990); Dvali and Vilenkin (2004) and a corresponding boost in the string density $\rho=\mu/L^{2}\propto P^{-1}$.. There are two main downsides to this approach. First, as shown in this paper, considering heavier string species in the cosmic superstring network is crucial to accurately model the resulting GW spectrum across its full frequency range. In particular, cosmic superstring networks have characteristic features like the presence of distinct peaks at high frequencies, or the domination of a different string species in different frequency ranges, and so modelling the GW signal through a single string network could lead to missed opportunities in potentially discovering these exciting features. However, it turns out that, within the NG15 frequency range, the fundamental string does in fact dominate and the one-string approximation is consistent with the data, as can be seen in Fig. 7 from the negligible contributions of the D and FD-strings. The second and main downside of the single-string approximation comes from the unclear physical interpretation of the parameter $P$. Although the one-string approximation can provide a best-fit likelihood and tension, the fundamental string parameters $g_{s}$ and $w$ are in fact encoded in $P$ and information about them can only be obtained through the full treatment outlined in this paper.
@@
Performing this analysis recreates the (cuspy) best-fit spectrum obtained by NANOGrav with the same fundamental string tension found in Afzal et al. (2023) and a corresponding best-fit string coupling and volume factor $g_{s}=0.06$ and $w=0.01$. The likelihood posteriors give 68% confidence intervals of $\log_{10}(G\mu_{1})^{NG}=-11.4^{+0.3}_{-0.3}$, and $2\sigma$ upper limits of $g_{s}^{NG}<0.54$ and $w^{NG}<0.82$. Note that the 68% interval on string tension is consistent with the values found by Afzal et al. (2023) (confidence interval $[-12.08,-11.50]$) and Figueroa et al. (2024) ($\log_{10}(G\mu_{1})^{NG}=-11.79^{+0.31}_{-0.19}$).

### v2

@@ Abstract
We study the stochastic gravitational wave background sourced by a network of cosmic superstrings and demonstrate that incorporating higher-mass string species, beyond the fundamental string, is crucial for accurately modeling the resulting gravitational wave spectrum across frequencies ranging from nanohertz to kilohertz. Using the multi-tension velocity-dependent one-scale model to evolve the cosmic superstring network, we perform several fits to the NANOGrav 15-year dataset and obtain expectation values for the fundamental string tension, string coupling and effective size of compact extra dimensions. We find that the cosmic superstring best-fits are comparable in likelihood to Supermassive Black Hole models, thought by many to be the leading candidate explanation of the signal. The implications of the best-fit spectra are discussed within the context of future gravitational wave experiments. We obtain expectation values for the fundamental string tension of $\log_{10}(G\mu_{1})=-11.4^{+0.3}_{-0.2}$($-11.5^{+0.3}_{-0.2}$) for gravitational waves originating from large cuspy (kinky) cosmic superstring loops and $\log_{10}(G\mu_{1})=-9.7^{+0.7}_{-0.7}$($-9.9^{+1.0}_{-0.5}$) for small cuspy (kinky) loops. We also place $2\sigma$ upper bounds on the string coupling, finding $g_{s}<0.7$ in all cases, and comment on the implication of our results for the effective size of the compact extra dimensions.
@@ I Introduction
Cosmic superstrings occur in certain string theory scenarios where fundamental strings (F-strings) and one-dimensional Dirichlet branes (D-strings) are stretched to cosmic scales by brane inflation Sarangi and Tye (2002); Jones et al. (2002); Majumdar and Christine-Davis (2002); Copeland et al. (2004). Unlike ordinary cosmic string networks, which generally consist of a single type of string with unit chopping probability when intersecting, cosmic superstrings have multiple different species (the F- and D-strings together with their bound states) and a reduced probability of intersecting. As they are string theory objects, the extra dimensions they experience along with their more complicated dynamics [^1] Note that non-trivial features like multiple species and reduced intercommutation probabilities can also arise in different contexts, notably in pure Yang-Mills theory Yamada and Yonekura (2022, 2023). significantly increases the difficulty of modelling the evolution of these networks. Nevertheless, remarkable strides have been made towards a consistent picture of cosmic superstring network evolution by modifying methods used to study cosmic strings Avgoustidis and Copeland (2010); Avgoustidis and Shellard (2008); Pourtsidou et al. (2011); Marfatia and Zhou (2024); Correia and Martins (2022); Blanco-Pillado et al. (2021). While there remain uncertainties in the detailed behaviour and interactions of cosmic superstring networks, we believe that we are at a point where detailed predictions can be made about the spectrum of stochastic gravitational waves arising from them Sousa and Avelino (2016).
@@ II.1 String Dynamics
where $\mathcal{P}_{i}\equiv\mathcal{P}_{ii}$ is the microphysical self-interaction probability for strings of type $i$. The prefactor $\tilde{c}$ ensures that for ${\mathcal{P}}_{i}=1$ one recovers the standard VOS value $\tilde{c}$ Martins and Shellard (2002) for the loop chopping efficiency, while the power $1/3$ is supported Avgoustidis and Shellard (2006) by numerical simulations of Nambu-Goto strings with a suppressed intercommutation probability, in both matter and radiation dominated cosmologies333 Earlier simulations of Nambu-Goto strings evolving in Minkowski spacetime may suggest a slightly stronger dependence Sakellariadou and Vilenkin (1990); Dvali and Vilenkin (2004) of $\tilde{c}$ on the intercommuting probability $P$, namely $\tilde{c}\propto P^{1/2}$.. Note that going from $\tilde{d}_{ij}$ in equation (5) to $\tilde{d}^{k}_{ij}=\tilde{d}_{ij}S^{k}_{ij}$ in equations (2-3) involves a kinematic factor $S^{k}_{ij}$ Avgoustidis and Copeland (2010); Pourtsidou et al. (2011) (also see Copeland et al. (2007); Salmi et al. (2008); Copeland et al. (2008)) quantifying the relative probability that the collision proceeds via the additive or subtractive channel Tye et al. (2005), i.e. $(p,q)+(p^{\prime},q^{\prime})\longrightarrow(p\pm p^{\prime},q\pm q^{\prime})$, producing a heavier or lighter string segment, respectively.
@@ III.1 Importance of heavier strings
We start by confirming that, as shown in Sousa and Avelino (2016), it is crucial to model the gravitational wave contributions of strings heavier than the fundamental string, as they give rise to additional structure in the spectrum. Since the cutoff frequency of GW emission of a given string species, $i$, is given by [^6] Recall we have the loop size given in terms of the time of chopping by $l_{i}=\alpha_{i}t$, and the fundamental frequency of the GW it emits is given by $f_{i}=2/l_{i}$. $f_{\rm min,i}=2/(\alpha_{i}t_{0})$, and since $\alpha_{i}$ can vary between string species, the cut-off frequency becomes species-dependent. If $\alpha_{i}$ differs between species, this leads to regions in frequency space where the GW emission has stopped for fundamental strings, but continues for heavier strings, as discussed earlier (cf. low frequency regions of Fig. 2 corresponding to small loops with $\alpha_{i}=2\Gamma G\mu_{i}$) for $g_{s}=0.04$ and $g_{s}=0.3$. The result is a multi-peaked spectrum at low frequencies – a clear signature of the GWs arising from a multi-tension network.
However, this is not the whole story. In addition to multiple peaks at low frequencies, there can also be areas in frequency space where GW emission from heavier strings dominates over emission from fundamental strings, even though the latter is not negligible (see leftmost plot of Fig. 3 corresponding to large loops with $\alpha_{L,i}=0.37$ for $g_{s}=0.04$). As the birth times of loops relevant to a given GW frequency (Eq. 11) depend on the tension of the string species (and loop density is cut off when $t<t_{b}$), the integral in Eq. (16) can be smaller for the fundamental strings than for heavier strings, despite the fact that the smaller correlation lengths result in many more loops of fundamental string overall. This effect is distinct from the appearance of multiple peaks (which is a direct consequence of having a hierarchy of loop sizes) and is present even when $\alpha_{i}$ is the same for all species.
In both cases, however, GW contributions from heavier string species are strongest for low values of the string coupling $g_{s}$ (for which the hierarchy in string tensions is higher) and show up at the low frequency part of the spectrum.
@@
Our results, summarised in Figs. 2 and 3 show that, in both the small loop and large loop limits, there are distinct signatures in the GW emission from heavier string species. These are more prominent for small $g_{s}$ as this gives rise to both a larger tension (leading to higher emission amplitude) and a larger loop size (leading to a lower frequency peak) for the heavy D and FD strings. These results are in agreement with reference Sousa and Avelino (2016), up to changes in the locations of the peaks caused by different choices/interpretations of the parameter $\alpha_{i}$, which in the case of reference Sousa and Avelino (2016) corresponds to $\alpha_{L,i}$ in our notation. Thus, the physical loop sizes differ (between this work and reference Sousa and Avelino (2016)) by a factor of twice the time-normalised correlation length, $2\xi_{i}$, in the small loop limit, and by a factor of 3.7 in the large loop limit. Other, smaller, differences arise from our more accurate modelling of the cosmic superstring network achieved by (a) including decay processes of heavier strings down to the three lightest species (as discussed above), and (b) taking into account the dependence of the VOS parameters $\tilde{c_{i}}$, $\tilde{d}^{k}_{ij}$ on the scaling velocities of each string species.
Figure 3: Plots showing the gravitational wave background spectra generated at different string coupling constants at string tension $G\mu_{1}=10^{-9}$ and loop size $\alpha_{L,i}=0.37$. The solid lines show the GWs from the 7-string model and the dashed lines from the 3-string model.
@@ III.2.2 Loop size, $\alpha_{i}$
One of the most important parameters in determining the shape and magnitude of the GW background spectrum is the size of string loops when they are chopped off from the string network. Motivated by the scaling behaviour of cosmic (super)string networks, it is commonly assumed that their size is a constant fraction of the string correlation length at the time of birth so that $l_{i}(t_{b})=\alpha_{i}t_{b}$, where $\alpha_{i}$ is a constant. The exact value of $\alpha_{i}$ is unknown. Throughout the years, investigations into cosmic strings have shown loops produced at sizes comparable to the string thickness Bevis et al. (2008); Vincent et al. (1997a), sizes around the gravitational backreaction scale, $\alpha_{i}=2\Gamma G\mu_{i}$ Bennett and Bouchet (1990); Albrecht and Turok (1985); Vincent et al. (1997b); Allen and Shellard (1990), and at scales much closer to the characteristic length of the string network Olum and Vanchurin (2007); Martins and Shellard (2006); Ringeval et al. (2007); Vanchurin et al. (2005, 2006). The most recent large-scale numerical simulation into the production of cosmic string loops finds a double-peaked spectrum of loop production, where around 90% of network energy goes into loops at the gravitational backreaction scale and the remaining 10% into large loops with $\alpha_{i}\sim 0.1$ Blanco-Pillado et al. (2014). Note in that paper the size of loops is expressed as a fraction of the horizon size during the radiation era, i.e $\alpha_{h}$, which they define as $\alpha$. This differs from the definition used in this work by a factor of 2.
Given the lack of a clear consensus, we will look at GWs emitted at the two scales found by Blanco-Pillado et al. (2014) as these will allow us to look at both the large and small loop physics of GW emission. The results of Blanco-Pillado et al. (2014) are obtained for radiation-era cosmic strings, which have a time-normalised correlation length of $\xi_{r}=0.27$, so that even in the large loop limit we have $\alpha_{i}<\xi_{r}$. As seen in Fig. 1, at small values of $g_{s}$, our correlation lengths can be as low as $\xi\sim 0.03$. As it would be unphysical to assume that, on average, loops are born with a size greater than the correlation length, we will instead fix $\alpha_{L,i}=0.37$. This recreates $\alpha_{i}=\alpha_{L,i}\xi_{i}=0.1$ for ordinary cosmic strings and ensures that loop sizes stay below the correlation length throughout the parameter space for cosmic superstrings. In the small loop limit, we will still set $\alpha_{i}=2\Gamma G\mu_{i}$ in agreement with Blanco-Pillado et al. (2014). Note that it is not clear that results derived from cosmic string simulations can be directly applied to cosmic superstrings and $\alpha_{i}$ can, in principle, take any other value. Nevertheless, we would expect these scales to be representative of the range of possible values.
@@ III.2.4 Size of the extra dimensions and energy of the loops
As discussed in section II, the volume parameter $w$ is the ratio between the minimum volume and effective volume occupied by the fundamental string in the compact extra dimensions. A small value of $w$ corresponds to a larger volume of the compact dimensions being available to the strings, while $w=1$ corresponds to compactification at the string scale (or to highly warped extra dimensions) so that volume effects from the extra dimensions become negligible. Note that the effective volume is only a lower bound to the total volume of the extra compact dimensions. As more of the extra dimensions become traversable to the strings, the larger the probability of the strings “missing” each other in the extra dimensions when intersecting. This leads to a decrease in intercommuting probability, corresponding to an increase in the number of string loops and a boost to the GW signal. Since the heavier strings oscillate less in the extra dimensions, the effective volume they explore is much smaller than for the fundamental strings and the boost in GW magnitude is much smaller.
Comparison spectra for $w=0.1$ and $w=1$ can be seen in Fig. 4. As varying $w$ has a strong impact on the resulting spectra and the value of $w$ is physically interesting, encoding information on the size/warping of the extra dimensions, we will treat this as a free parameter.
@@ IV Results and Discussion
Cuspy big loops $47.9$ $-11.4^{+0.3}_{-0.2}$ $<0.69$ $<0.90$
-11.8 0.22 0.01
Kinky big loops $47.8$ $-11.5^{+0.3}_{-0.2}$ $<0.70$ $<0.91$
-11.9 0.26 0.01
@@ IV.3 Comparison to NANOGrav fit
Figure 8: Likelihood posteriors obtained from fitting the model considering $10\%$ large loops and $90\%$ non-contributing small loops to the NG15 data. The solid (dashed) lines correspond to cuspy (kinky) loops.
@@
where $\Omega_{\rm CS}$ is the spectrum obtained from cosmic strings at a given tension and $P$ is treated as a constant free parameter [^7] The scaling $\Omega_{\rm SUPER}(f)\propto P^{-1}$ can be understood in terms of the VOS model as taking the loop chopping efficiency $\tilde{c}\propto P^{1/2}$ leading to a scaling $L\propto P^{1/2}$ in the correlation length Sakellariadou and Vilenkin (1990); Dvali and Vilenkin (2004) and a corresponding boost in the string density $\rho=\mu/L^{2}\propto P^{-1}$.. There are two main downsides to this approach. First, as shown in this paper, considering heavier string species in the cosmic superstring network is crucial to accurately model the resulting GW spectrum across its full frequency range. In particular, cosmic superstring networks have characteristic features (the presence of multiple peaks and/or domination of heavy string species at low frequencies) and so modelling the GW signal through a single string network could lead to missed opportunities in potentially discovering these exciting features. However, it turns out that, within the NG15 frequency range, the fundamental string does in fact dominate and the one-string approximation is consistent with the data, as can be seen in Fig. 7 from the negligible contributions of the D and FD-strings. The second and main downside of the single-string approximation comes from the unclear physical interpretation of the parameter $P$. Although the one-string approximation can provide a best-fit likelihood and tension, the fundamental string parameters $g_{s}$ and $w$ are in fact encoded in $P$ and information about them can only be obtained through the full treatment outlined in this paper.
@@
Performing this analysis gives cuspy (kinky) likelihood posteriors with 68% confidence intervals of $\log_{10}(G\mu_{1})^{NG}=-11.0^{+0.4}_{-0.2}$($-11.1^{+0.4}_{-0.3}$), and $2\sigma$ upper limits of $g_{s}^{NG}<0.59$($<0.61$) and $w^{NG}<0.84$($<0.84$). The posterior likelihood marginals obtained can be seen if Fig. 8. Note that the 68% interval on the string tension is significantly different to that found by Afzal et al. (2023) (confidence interval $[-12.08,-11.50]$) and Figueroa et al. (2024) ($\log_{10}(G\mu_{1})^{NG}=-11.79^{+0.31}_{-0.19}$), as we are using smaller loop sizes in accordance with our more accurate modelling of the string correlation lengths.
This discrepancy highlights another shortcoming of the $1/P$ modelling approach: its failure to accurately account for variations in correlation length. By calculating the ordinary cosmic string GW signal and then rescaling it by a factor of $1/P$, this approach, in effect, takes the correlation length to scale as [^8] Also note that, while the scaling $L\propto\sqrt{P}$ is consistent with flat space simulations Sakellariadou and Vilenkin (1990) and early theoretical expectations Dvali and Vilenkin (2004), string evolution simulations in an expanding universe find a weaker dependence, $L\propto P^{1/3}$, in both the matter and radiation eras Avgoustidis and Shellard (2006). $L\propto\sqrt{P}$ without a corresponding rescaling of the loop size, which is fixed to $\alpha=0.1$. This is relevant in the cosmic superstring case, as discussed in section III.2.2, where fixing $\alpha_{i}=0.1$ can lead to loops with lengths multiple times longer than the correlation length of each string species. In fact, when we impose $\alpha_{i}=0.1$ in our analysis (along with the $0.1$ factor in the loop number density) we recover tension bounds close to those in Afzal et al. (2023); Figueroa et al. (2024), and reproduce NANOGrav’s best-fit spectrum exactly. The best-fit tension of $log_{10}(G\mu_{1})=-11.94$, however, corresponds to $g_{s}=0.06$ and $w=0.01$, implying a fundamental string correlation length of $\mathcal{O}(0.01)$, as shown in Fig. 1. At such small correlation lengths, the loop sizes used to obtain this spectrum become unphysically large.